OpenStudy (anonymous):

Hi--- I don't mean to ask for much but my SAT is tomorrow, and I just want to be perfectly prepared for it!! So, I was wondering if you guys could help me understand a few questions on this SAT preparation?

OpenStudy (anonymous):

A fruit salad is made from pineapples, pears, and peaches mixed in the ratio of 2 to 3 to 5, respectively, by weight. What fraction of the mixture by weight is pineapple? (A) 1/5 (B) 3/10 (C) 2/5 (D) 1/2 (E) 2/3

OpenStudy (anonymous):

I have no idea how to do ratios, I tried on Khan and it wasn't exactly what I was hoping/looking for...

hero (hero):

Basically, for this one, you can assume you have 2 pineapples, 3 pears and 5 peaches and a total of 10 pieces of fruit. So the ratio of pineapples to total fruit is 2/10 or 1/5 So therefore, the pineapples will carry 1/5 of the mixture by weight.

OpenStudy (anonymous):

Oh-- Well, you made that seem easy! xDD Alright, so if we were to do... Something like, A pet store has various kinds of animals from fish, dogs, and cats mixed in the ratio of 8 to 12 to 16, respectively, by weight. What fraction of the mixture by kinds is fish? 8+12+16= 36 8/36?

hero (hero):

I know it's really a mixture ratio in weight, but it's just more convenient to describe it in terms of pieces of fruit.

hero (hero):

Don't forget to reduce the fraction 8/36 = 2/9

OpenStudy (anonymous):

By dividing it by the GCF?

hero (hero):

Yes, you should know how to reduce fractions by now. You're about to take the SAT.

OpenStudy (anonymous):

 The graphs of the functions f and g in the interval from x=-2 to x=2 are shown above. Which of the following could express g in terms of f? (A) g(x)=f(x+1) (B) g(x)=f(x)+1) (C) g(x)=f(x+1)+1 (D) g(x)=f(x-1) (E) g(x)=f(x)-1